OpenStudy (anonymous):

Find the equation of the tangent line at (-1,5) on the circle x^2+y^2+2x-9=0

OpenStudy (phi):

Do you know how to do implicit differentiation?

OpenStudy (phi):

take the derivative of each term with respect to x use these rules: \[ \frac{d}{dx} x = \frac{dx}{dx} =1 \\ \frac{d}{dx} y= \frac{dy}{dx} \]

OpenStudy (anonymous):

no :(

OpenStudy (phi):

can you do \[ \frac{d}{dx} x^2 \]?

OpenStudy (anonymous):

substitute it to (-1,5)?

OpenStudy (phi):

no, using the (-1,5) comes at the end. First we have to find the slope of the tangent line to the curve. that means we find dy/dx (which by definition is the slope of the tangent line) and *that* means use differential calculus you need to first take the derivative of your function. The first step is, can you find the derivative of x^2 with respect to x?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[******x ^{2}+y^{2}+4x-9=0 ******\]

OpenStudy (anonymous):

@phi (x+2)^2?

OpenStudy (phi):

it looks like you completed the square, which is a way to put the equation into the standard form of a circle. That is useful for finding the center of the circle. But, it is not needed for this problem.

OpenStudy (anonymous):

this might actually not be a calculus problem i have seen it done without using derivatives it is a pain, derivatives make it much easier but if this is not calc class then this will make no sense

OpenStudy (phi):

oh! @SeniorHigh what course is this ??

OpenStudy (anonymous):

introduction to analytical geometry&calculus

OpenStudy (phi):

Have you learned (or were you supposed to have learned) how to take a derivative?

OpenStudy (anonymous):

we didn't discuss it yet.. still on the circles (CONICS)

OpenStudy (anonymous):

thought so damn

OpenStudy (phi):

In that case, we should plot the circle and the point (-1,5) because we are going (painful) geometry

OpenStudy (anonymous):

i think there is some method arguing that when you solve the simultaneous equation the discriminant must be zero if they are to touch let me see if i can remember it

OpenStudy (anonymous):

sir what is the center and the slope of radius?

OpenStudy (anonymous):

oh no it is not that bad at all

OpenStudy (phi):

I can see how to do it using ratios of similar triangles. 1) the radius will form a right angle with the tangent line

OpenStudy (anonymous):

1) find the center of the circle by completing the square 2) find the slope of the line from the center to the point \((-1,5)\) 3) the line will be perpendicular, take the negative recprocal 4) use the point slope formula

OpenStudy (anonymous):

@phi seem reasonable?

OpenStudy (anonymous):

we can walk through it if you like, i am fairly certain that this will work

OpenStudy (phi):

(-1,5) is not on the circle

OpenStudy (anonymous):

oh damn

OpenStudy (anonymous):

i must have read the problem wrong, thought it said "one the circle"

OpenStudy (phi):

but we do have similar right triangles. we know length 5, "height" (i.e. radius) sqr(10), and hypotenuse sqr(15)

OpenStudy (anonymous):

do i need to perform the completing the square in \[x ^{2+y ^{2}+4x-9=0}\]?

OpenStudy (phi):

Yes, to do this problem you need to at least know its center and radius can you complete the square on the x terms?

OpenStudy (anonymous):

\(4x\) or \(2x\) ?

OpenStudy (anonymous):

4x @satellite73

OpenStudy (anonymous):

that shoots all my calculations to hell

OpenStudy (anonymous):

\[(x+2)^{2}+(y-3)^{2}\] is this right?

OpenStudy (anonymous):

you sure? in the question you wrote \[x^2+y^2+2x-9=0\]

OpenStudy (phi):

Gack! the 4x makes it harder! and, there is now y to the first power, so you do not complete the square on the y part of the problem

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

typo @satellite73

OpenStudy (anonymous):

i know! that's why i hate math T.T

OpenStudy (anonymous):

so far i am at \[y^2=13-(x+2)^2\\ y=m(x+1)-5\]

OpenStudy (anonymous):

now i will try to square the bottom one try to solve for \(x\) but only set the discriminate equal to zero because there should only be one solution, then solve that equation for \(m\)

OpenStudy (phi):

there will be two tangent lines

OpenStudy (anonymous):

yes, when i solve for \(m\) it will be a quadratic, should get two solutions

OpenStudy (phi):

meanwhile, did @SeniorHigh find the equation of the circle ?

OpenStudy (anonymous):

this would be a ton easier if it was \(x^2+y^2+2x-9=0\)

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

but according to my classmate x+5-24=0

OpenStudy (phi):

Remember we are trying to form an equation of the form \[ (x-a)^2 + (y^2 - b^2) = r^2 \] Let's complete the square on the x part \[ x^2+4x+y^2−9=0 \\ x^2+4x+y^2=9 \] first, notice there is no y term. we don't have to do anything for the y^2 term next, notice the 4x term. we divide the 4 by 2. 4/2= 2. square it. 2*2= 4 that means we want a +4. add +4 to both sides \[ x^2+4x+4 +y^2=9+4 \] the x^2+4x+4 = (x+2)^2 (multiply out to show it's true) \[ (x+2)^2 +y^2= 13\]

OpenStudy (phi):

based on that equation, can you identify the center of the circle and its radius?

OpenStudy (anonymous):

yes but can you explain how to do it?

OpenStudy (phi):

yes what?

OpenStudy (phi):

I think you mean no, can you explain it? match the equations \[ (x+2)^2 +y^2= 13 \\ (x-a)^2 + (y^2 - b^2) = r^2 \] the center will be (a,b) we have x+2 but that does not match x-a. but we can write it as (x- -2) we have just y^2, but we could write it (y-0)^2. we don't have an r^2, but we can write \( 13= (\sqrt{13})^2 \)

OpenStudy (phi):

so match these equations \[ (x - -2)^2 +(y-0)^2 = (\sqrt{13})^2 \\ (x-a)^2 + (y-b)^2 = r^2 \] what is a and b? (the center will be (a,b) ) what is r ?

OpenStudy (anonymous):

center is -2,0 and r is 5

OpenStudy (phi):

yes, -2,0 for the center. r is not 5. you match up r^2 with (sqr(13))^2 in other words, r matches sqr(13)

OpenStudy (anonymous):

oohh.. I get it now.. thank you :)

OpenStudy (phi):

you should plot (or sketch) the problem. Here it is

OpenStudy (phi):

we can find the lengths of the legs of the right triangle. 1) the length of the "short leg" is the radius. we know that is sqr(13) 2) what is the distance from the center to point A ? this is the hypotenuse 3) from 1) and 2) we can find the 3rd leg using Pythagoras

OpenStudy (phi):

you should find the "sides" are sqr(13) and the hypotenuse is sqr(26) in other words a 45-45-90 triangle I can see how to use trig (tangent of sum of two angles) to find the slope of the tangent line. There are probably other approaches

OpenStudy (phi):

Here is a picture

OpenStudy (phi):

the last step to find the equation of the line is y-5 = m(x - -1) or y-5= m(x+1) where m is the slope, or the perpendicular -1/m (for the other tangent line)